OpenStudy (anonymous):

How many numbers between 0 and 1000 are divisible by 5?

hero (hero):

Hint: It's an odd number. It's very close, but not 200

hero (hero):

Key word is "between"

OpenStudy (anonymous):

199

OpenStudy (anonymous):

actually the 1st divisible number is 5 , and last is NOT 1000; but 1000-5 = 995 since u said BETWEEN. So we have an arithmetic progression. 5 , 10 , 15....995. with 1st term A = 5 , common difference D = 5 and Last term L = 995. L = A + (N-1)D. u can solve hence for N.

OpenStudy (anonymous):

195?

OpenStudy (anonymous):

I dont think it would be 195. well uh, 995 = 5 + (n-1)5 990 = (n-1)5. divide both sides of equation by 5 198 = n-1.

OpenStudy (anonymous):

199?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes it is..:)